E 3 points The indicial equation of: x2y + xy + (x2 - 4)y = 0 is A)r? + 4 = 0 B) r2 - 4 = 0 C)r? - 4r = 0 D) r2 – 4r + 4 =
Answer #1

Ang x? tay + (x²A) y = 0 Here P(x)= x² Q(x)=x R(x) = x²-4 1 Ky ko lim (x-xo) a (x) P(x) (x-x.)? R (2) lim % (Ro) . N No P(when lim 1 Por lim de o e P (2) x 7 eso - lim ro x² x² 22 -1 Por Loolim x² R(x) P(a) lim x² (x² - 4) xo xt o 82x² _ À = O- A (1 20 = lim - 4 2o -A when she 2 The indicial equation is in r(8-1) + por +20 Po = 1, 2o – A the formA - 0 r(8-1) + 1 (0) 8²_r tr - 4 = 0 82_A=0 Hence indicial equation 8²4 option is right answer. 8 4 - O 13